OpenStudy (anonymous):

Which of the following forms a proportion with 35/21?

OpenStudy (anonymous):

\[\frac{ 25}{ 11 } or \frac{ 6 }{ 10 } or \frac{ 21 }{ 7 }or \frac{ 20 }{ 12 }\]

OpenStudy (anonymous):

Medal to best answer

OpenStudy (abb0t):

*best answer*

OpenStudy (anonymous):

?

OpenStudy (uri):

@keysf102 Did you try doing it yourself? or you want *best answer?*

OpenStudy (anonymous):

i found 20/12

OpenStudy (uri):

Show work.

OpenStudy (anonymous):

\[\frac{ 35 }{ 21 }=\frac{ 20 }{ 12 }\]

OpenStudy (uri):

Really..? How?